### v2

# In Search of Truth: An Interrogation Approach to Hallucination Detection

Abstract
Despite the many advances of Large Language Models (LLMs) and their unprecedented rapid evolution, their impact and integration into every facet of our daily lives is limited due to various reasons. One critical factor hindering their widespread adoption is the occurrence of hallucinations, where LLMs invent answers that sound realistic, yet drift away from factual truth. In this paper, we present a novel method for detecting hallucinations in large language models, which tackles a critical issue in the adoption of these models in various real-world scenarios. Through extensive evaluations across multiple datasets and LLMs, including Llama-2, we study the hallucination levels of various recent LLMs and demonstrate the effectiveness of our method to automatically detect them. Notably, we observe up to 62% hallucinations for Llama-2 in a specific experiment, where our method achieves a Balanced Accuracy (B-ACC) of 87%, all without relying on external knowledge  [^1] Our code, datasets, and task prompts can be found here. .
**footnotetext: Denotes equal contribution.
1 Introduction
Human studies have shown that people tend to be inconsistent when they are not telling the truth Brewer et al. (1999). As such, a common interrogation technique consists of repeated interviews that attempt to challenge the interviewer’s consistency in order to assess its credibility Granhag and Strömwall (2001). Truth tellers’ answers are well-grounded in their memory, hence, inconsistencies in the respondent’s answers are a strong indication of her not telling the truth Brewer et al. (1999); Dianiska and Meissner (2023). Inspired by these studies, we present a novel method for hallucination detection in LLMs. Our approach, which we call InterrogateLLM , employs a systematic evaluation of model-generated responses for potential hallucinations by repeating the process of reconstructing a query from its generated answer.
Repeated interviews are a very common and effective verification technique for human interrogations, however, it is not foolproof. In some cases, respondents manage to provide repeated false states that are consistent, while in other cases, truth-tellers may provide inconsistent responses due to memory errors Bartlett (1995). In a similar fashion, our method is not flawless; it represents an additional step towards addressing the yet unsolved problem of hallucination detection. Nevertheless, similar to the use of consistency tests in humans, commonly employed for their effectiveness, our method also demonstrates high efficacy.
In recent years, the emergence of LLMs such as GPT-3 Brown et al. (2020b), PaLM Chowdhery et al. (2022), and LLama Touvron et al. (2023a, b) has revolutionized natural language processing. These models enable machines to understand and generate human-like text with unprecedented fluency and coherence. These models, trained on vast amounts of text data, have demonstrated remarkable capabilities in various applications, from automated content generation to virtual assistants, and beyond. However, their remarkable performance comes with a set of challenges and concerns that need to be addressed for their responsible and effective use. A major concern is the phenomenon of hallucination, whereby these language models generate misleading, potentially harmful, or erroneous text. Hallucination can be characterized by the presence of false information in the output generated by the language model that lacks a factual basis. There are significant challenges associated with the deployment of large language models in real-world applications, especially in those involving critical information or decision-making processes.
Detecting and minimizing hallucinations in LLMs is crucial for ensuring their trustworthiness and reliability, especially in contexts where these models play a pivotal role in communication and decision-making. Existing methods for evaluating model-generated text often rely on surface-level metrics such as fluency and coherence, which may not effectively capture the underlying issue of hallucinations. Therefore, there is a pressing need for a systematic and effective method to detect and mitigate hallucinations in the outputs of these models. Despite its significance, addressing this challenge remains an open issue Ji et al. (2023).
Our method, InterrogateLLM , operates on the premise that language models exhibiting hallucinations produce inconsistent and incorrect responses to subsequent queries based on the hallucinated information. To identify hallucination in a generated answer, our approach involves prompting the model multiple times to reconstruct the input query using the generated answer. Subsequently, InterrogateLLM quantifies the inconsistency level between the original query and the reconstructed queries. By leveraging the observed inconsistencies, our approach effectively identifies potential instances of hallucination. When a large-language model generates a hallucination, it struggles to consistently reconstruct the original query, leading to variations in responses. This interrogation strategy serves as the cornerstone of our approach for detecting hallucinations in generated answers.
The contributions of our paper are outlined as follows: (1) introduction of the InterrogateLLM method designed for detecting hallucinations in textual answers generated by LLMs. (2) we propose an innovative evaluation approach specifically tailored to the task of hallucination detection, leveraging three datasets associated with our proposed text generation tasks. (3) we investigate the hallucination levels exhibited by recent LLMs, including Llama2, shedding light on their fidelity levels. (4) we present comprehensive performance reports on InterrogateLLM and its variants, conducting a thorough comparison with alternative methods through extensive evaluations.
Figure 1: An illustration of the InterrogateLLM method. (1) A few-shot prompt and a query are fed into $F_{LLM}$, which generates an answer. (2) The shots in the prompt are then reversed, forming a sequence of answer-question pairs, with the generated answer constructed on top. The $B_{LLM}$ is then used to generate $K$ queries that correspond to the generated answer. Ideally, the generated queries should recover the original query from the forward phase. (3) The set of recovered questions is then embedded by a language model and compared with the original question, producing a final score that determines whether the generated answer suffers from hallucination.
2 Related Work
Hallucinations have been explored in various natural language generation tasks, including translation, summarization Kryscinski et al. (2020); Maynez et al. (2020), dialogue generation Shuster et al. (2021), and question-answering Lin et al. (2022). This is well-documented in a recent comprehensive survey conducted by Ji et al. (2023), which provides an insightful overview of hallucinations in diverse natural language generation contexts.
In Liu et al. (2022), the authors presented a token-level reference-free hallucination detection task along with an additional dataset designed for hallucination detection in free-form text. This dataset consists of textual passages with perturbations, and the objective is to determine whether the entire passage exhibits hallucinations. It is crucial to emphasize that our task differs from their setup, as we specifically address hallucination detection within few-shot prompts involving query-answer sequences.
To address inconsistencies in generated text, SelfCheckGPT, introduced by Manakul et al. (2023b), leverages multiple stochastic samples generated by LLMs using the same query. SelfCheckGPT evaluates the coherence between the response and the stochastic samples by querying the same LLM multiple times. Specifically, it employs an additional prompt that includes a stochastic sample and a sentence from the generated text and predicts whether the sentence is supported by the stochastic sample. The methodology of SelfCheckGPT encompasses two distinct approaches, one is based on BERTScore Fu et al. (2023), and the other employs a multiple-choice question answering and generation approach (MQAG) Manakul et al. (2023a).
In recent research, Azaria and Mitchell (2023) proposed a method employing a multi-layer perceptron classifier that uses hidden representations from language models to predict sentence truthfulness. However, this approach necessitates labeled data for supervised training and access to the internal states of the language model, which may not always be readily available.
In Kadavath et al. (2022), the authors present a self-evaluation technique where models are trained to predict their knowledge of the answer to any given free-form question. This approach entails prompting the language model to internally assess the accuracy of its previous predictions, including estimating the likelihood that its generated response or answer is correct. It is worth noting that this method requires labeled data for model training, making it a supervised task, which differs from our settings.
3 Problem setup
We assume a source domain of textual queries and a target domain of textual answers. A few-shot prompt [^2] While our approach assumes a provided few-shot prompt, it stays adaptable to many zero-shot tasks where the creation of few-shot prompts is feasible. Brown et al. (2020b), a corresponding query $Q$ and a LLM denoted by $F_{LLM}$, are provided. The query is constructed on top of the prompt and fed into the LLM to generate an answer to the query. Our task is to detect whether the generated answer suffers from hallucinations.
The few-shot prompt is constructed as a sequence of query-answer pairs. The pairs are denoted by $\left\{\left(q_{i},a_{i}\right)\right\}_{i=1}^{n}$, where $q_{i}$ represents a query and $a_{i}$ its corresponding answer. The prompt can be expressed as follows:
$\mathcal{P}_{forward}=q_{1},a_{1},\ldots,q_{n},a_{n}$ (1)
The $F_{LLM}$ is queried with the concatenation of the query $Q$ on top of the prompt $\mathcal{P}_{forward}$, which retrieves a generated answer denoted by $A^{*}$, signifying the response to the query $Q$. In other words, the prompt $\mathcal{P}_{forward}$ and the query $Q$ are fed into the LLM as follows:
$A^{*}=F_{LLM}([\mathcal{P}_{forward};Q])$ (2)
Our task is to determine whether the generated answer $A^{*}$ exhibits hallucinatory content.
4 The InterrogateLLM method
In our approach, we introduce a backward process for reconstructing the original query $Q$ from the generated answer $A^{*}$. We create a new prompt by reversing the given prompt $\mathcal{P}_{forward}$. The reversed prompt rearranges the order of the query-answer pairs to pairs of answer-query. The reversed prompt, denoted as $\mathcal{P}_{backward}$, can be expressed as follows:
$\mathcal{P}_{backward}=a_{1},q_{1},\ldots,a_{n},q_{n}$ (3)
The generated answer $A^{*}$ is then concatenated to the end of the reversed prompt $\mathcal{P}_{backward}$, and the entire sequence is passed either by the same LLM defined above, by a different LLM, or by an ensemble of LLMs. For ease of reference and clarity, we collectively refer to the LLMs involved in this step as $B_{LLM}$. In other words, in this process, we map the generated answer to the source domain, by querying one or more LLMs, each trying to reconstruct the original query $Q$. By denoting the set of reconstructed queries as $Q^{*}$, this “backward” process can be expressed as:
$Q^{*}=B_{LLM}([\mathcal{P}_{backward};A^{*}])$ (4)
Note that the size of $Q^{*}$ depends on the number of LLMs used in $B_{LLM}$.
The motivation for employing a backward process is to reconstruct the original query $Q$ based on the generated answer $A^{*}$. If the initial LLM suffers from hallucinations during the generation of $A^{*}$, then $A^{*}$ may drift from the correct answer to $Q$. Consequently, a backward process operating on $A^{*}$ is prone to deviating from $Q$ on the way back. In other words, in the case of hallucination in $A^{*}$, the set of reconstructed queries $Q^{*}$ is likely to diverge from the original query $Q$.
Input: $Q$, $F_{LLM}$, $B_{LLM}$, $M_{embedding}$, $\left\{\left(q^{(i)},a^{(i)}\right)\right\}_{i=1}^{n}$, $\tau$
Output: True if the generated answer is an hallucination, False otherwise
1 Step 1: Query Forward Pass $\mathcal{P}_{forward}=q_{1},a_{1},\ldots,q_{n},a_{n}$  $A^{*}=F_{LLM}([\mathcal{P}_{forward};Q])$ Step 2: Query Reconstruction $\mathcal{P}_{backward}=a_{1},q_{1},\ldots,a_{n},q_{n}$  $Q^{*}=\{\}$ for $i=1$ to $K$ do
2       Substep 1: Reverse Pass $Q^{*}=Q^{*}\cup B_{LLM}([\mathcal{P}_{backward};A^{*}])$
[…]
4       return False
5else
6       return True
Algorithm 1 Hallucination Detection
In InterrogateLLM , the backward process is repeated multiple times ($K$ times for each model in $B_{LLM}$, see Sec. 5.2 for more details). Therefore,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}|Q^{*}|=K*|B_{LLM}|}$
To determine if $A^{*}$ suffers from hallucination, a language embedding model is utilized to assess the similarity between the set of reconstructed queries $Q^{*}$ and the original query $Q$. Both the generated queries and the original query are transformed into vectors within the same embedding space. For a given embedding model $M_{embedding}:text\rightarrow\mathbb{R}^{D}$, which generates $D$-dimensional vectors from the input text, the embedding vector for the original query $Q$ is denoted as $Q_{vec}=M_{embedding}(Q)$. Similarly, the embedding vectors for the generated queries are denoted by:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Q^{*}_{vec}=\{M_{embedding}(q^{*}):\forall q^{*}\in Q^{*}\}}$
Subsequently, the cosine similarity between the embedding vectors of the predicted queries $Q^{*}_{vec}$ and the original query $Q_{vec}$ is calculated as follows:
$\text{sim}(Q,Q^{*})=\underset{\forall q^{*}_{\text{vec}}\in Q^{*}_{\text{vec}}}{\text{AVG}}\left(\left[C(Q_{\text{vec}},q^{*}_{\text{vec}})\right]\right)$ (5)
Here, $C$ represents the cosine similarity function:
$C(u,v)=\left(\frac{u\cdot v}{\|u\|\|v\|}\right)$ (6)
for $u,v\in\mathbb{R}^{D}$, where $D$ is the dimension of the vectors. In other words, the cosine similarity is calculated for each $q^{*}_{\text{vec}}$ in the set $Q^{*}_{\text{vec}}$, and the results are then averaged to obtain the final similarity score.
Finally, InterrogateLLM predicts hallucinations if the similarity score exceeds a predetermined threshold $\tau$. In essence, when the reconstructed queries exhibit a significant divergence from the original query, InterrogateLLM signifies that there is a potential hallucination in $A^{*}$. More details about the selection of $\tau$ can be found in Sec. 5.2. The InterrogateLLM method is illustrated in Fig. 1, and outlined in Alg. 1.
4.1 Variable temperatures
We introduce an exploratory extension into InterrogateLLM , exploring the impact of diverse temperature values on the accuracy of the detections. In standard LLMs, the temperature parameter influences the likelihood of selecting the next token during the answer generation process. A higher temperature (e.g., 1.0) makes the output more creative and diverse, while a lower temperature (e.g., 0.2) makes the output more focused and deterministic. Specifically, the temperature is applied through a softmax function that transforms a vector into a probability distribution. In text generation, the softmax function is applied to the model’s logit vector, which corresponds to the supported tokens in the vocabulary. The softmax operation can be written as follows:
$P_{i}=\frac{e^{{z_{i}}/{T}}}{\sum_{j=1}^{N}e^{{z_{j}}/{T}}}$ (7)
Where $P_{i}$ is the probability of selecting the $i$-th token in the vocabulary, $z$ is the logit vector, $T$ is the temperature parameter and $N$ is the number of tokens in the vocabulary. When $T$ is high (low), the exponential function ${e^{{z_{i}}/{T}}}$ is less (more) sensitive to small differences in the logit values, making the probabilities more diverse (focused).
As complementary experimental explorations, we examine the influence of temperature values on InterrogateLLM during the backward process, which is iterated $K$ times. By introducing dynamic temperature adjustments, our goal is to study the method’s accuracy when employed with a range of backward processes exhibiting diverse creativity levels. To this end, we set the temperature for each backward process as follows:
$T_{i}=T_{0}+\frac{1.0-T_{0}}{K}\cdot i$ (8)
where $T_{i}$ represents the temperature for the $i$-th backward pass ($0\leq i<K$), and $T_{0}$ is the model default temperature (see Sec. 5.2 for more details).
This temperature scheduling allows for facilitating a controlled ascent in temperatures across the multiple backward processes, promoting enhanced exploration in the space of reconstructed queries. The details and results of this additional study are reported in the experiments, Sec. 5.6.
Hallucination Rate
$F_{LLM}$
Movies
Books
GCI
GPT3
37%
38%
0%
Llama-2 (7B)
87%
66%
25%
Llama-2 (13B)
72%
58%
60%
Table 1: Hallucination rates for each dataset and $F_{LLM}$.
Movies Books GCI
$F_{LLM}$ Method   AUC    B-ACC     AUC    B-ACC     AUC    B-ACC
GPT3 $InterrogateLLM$ $B_{LLM}$ GPT3 0.817 0.739 0.709 0.673 - 0.994
Llama-2 (7B) 0.751 0.639 0.646 0.616 - 0.983
Llama-2 (13B) 0.789 0.695 0.684 0.640 - 0.983
Ensemble 0.818 0.699 0.710 0.656 - 0.983
SBERT-cosine 0.616 0.500 0.534 0.500 - 0.550
ADA-cosine 0.709 0.500 0.530 0.500 - 0.591
Llama-2 (7B) $InterrogateLLM$ $B_{LLM}$ GPT3 0.824 0.786 0.828 0.787 0.965 0.952
Llama-2 (7B) 0.823 0.750 0.761 0.707 0.959 0.958
Llama-2 (13B) 0.828 0.775 0.795 0.734 0.969 0.960
Ensemble 0.874 0.813 0.822 0.761 0.951 0.948
SBERT-cosine 0.586 0.516 0.552 0.486 0.957 0.548
ADA-cosine 0.770 0.501 0.641 0.499 0.950 0.820
Llama-2 (13B) $InterrogateLLM$ $B_{LLM}$ GPT3 0.806 0.753 0.804 0.754 0.989 0.982
Llama-2 (7B) 0.788 0.706 0.742 0.697 1.000 1.000
Llama-2 (13B) 0.801 0.746 0.771 0.709 0.995 0.991
Ensemble 0.842 0.773 0.807 0.733 0.992 0.964
SBERT-cosine 0.539 0.505 0.573 0.497 0.955 0.546
ADA-cosine 0.728 0.500 0.600 0.500 0.966 0.852
Table 2: Hallucination detection results for all models and datasets. InterrogateLLM is reported with $K=5$ and variable temperature values. For each dataset and $F_{LLM}$, we compare InterrogateLLM and its variants to all other baselines. As GPT3 does not suffer from hallucinations on the GCI dataset, only the B-ACC metric is reported.
5 Experiments
To assess the efficacy of InterrogateLLM in detecting hallucinations, and due to the absence of prior datasets for hallucination detection in few-shot prompt settings, we adapted three public datasets. For each dataset, we designed a text generation task along with a verification process to ensure the accuracy of the generated answers. The verification is implemented by employing simple heuristic functions that exploit additional information that is present in the datasets. During the evaluation of hallucination detection methods, the detection predictions are compared against the verification results. Importantly, the InterrogateLLM method operates independently of any external knowledge, making it versatile and applicable to a broad spectrum of tasks.
5.1 Datasets and Tasks
A comprehensive experimental evaluation was conducted using three different datasets to thoroughly evaluate our hallucination detection method across various domains. All three datasets provide a multifaceted evaluation of our technique, revealing its versatility across various types of information and content and allowing us to test the robustness of our hallucination detection method across a wide range of datasets and domains.
5.1.1 The Movies Dataset
The Movies Dataset [^3] The Movies Dataset is a collection of movie-related data that is publicly available for analysis and research. The dataset contains a variety of details about movies that were released before July 2017. The dataset includes 26 million ratings and 750,000 tag applications for all 45,000 movies provided by 270,000 users.
A subset of 3000 samples with movie titles and release years associated with the movie cast was sampled from the Movies dataset. The task is to predict the cast of a movie based on the movie’s name and release year. The few-shot prompt contains a few examples mapping a movie’s name and release year to its cast. The prompt is in the following format: "Query: What actors played in the $x$ movie $y$?" where $x$ is the release year and $y$ is the movie name. Cast members’ full names are expected in answers, and ground truth labels use Intersection Over Union (IOU) scores, considering any IOU score below 80% as a hallucination.
5.1.2 Books Dataset
The second dataset ("books dataset [^4] Books Dataset)" is derived from Amazon and includes over 200,000 literary books. This public dataset provides an overview of diverse literary books available on the Amazon platform. Each record includes details like book title, authors, publishers, and publication year.
We sampled a subset of 3,000 samples, including titles, dates, authors, and publishers. The task is to predict the author and publication year based on the book title. The prompts are structured as "Who is the author of the book $x$, what year was it published?", where $x$ is the book title. The ground truth is established by checking for a match between the elements (author name, release year) in the answer.
5.1.3 Global Country Information (GCI)
The “Global Country Information” [^5] GCI Dataset (GCI) is a public dataset containing information on 195 countries. Detailed information about each country is provided, including its name, land area, capital or major city, GDP, and more. This dataset offers a comprehensive representation of global country information. In the GCI dataset, we concentrate on country and capital pairs. The task involves determining a country’s capital by asking, "What is the capital of $x$?"
Samples from the above three datasets can be found in the supplementary Sec. B. The prompts used in each datasets and the reversed prompts created by InterrogateLLM can be found in the code.
5.2 Implementation details
We set $K=5$ and $\tau=0.91$ across all experiments. Maintaining a relatively small value for $K$ facilitates rapid benchmarking of various models on datasets in our evaluations, that encompass tens of thousands of generated answers. The hyperparameter $\tau$ was determined through an analysis of ada002 embeddings on a third-party dataset. This involved embedding both similar and dissimilar sentence pairs within the QQP dataset Chen et al. (2018) and selecting the optimal threshold that effectively distinguished between the two distributions. The initial temperature $T_{0}$ was set to the default temperature of each of the evaluated LLMs, specifically $0.6$ for GPT3 and both Llama-2 models. The embedding model used in InterrogateLLM leverages the latest OpenAI’s model, ada002666https://platform.openai.com/docs/guides/embeddings/use-cases.
In our experiments, we utilized a single A100 GPU. The execution time for a single application of InterrogateLLM , with the full method, leveraging an ensemble of three models, takes up to 2 seconds. Consequently, benchmarking InterrogateLLM across the three datasets takes up to $\sim 3.44$ hours. Further insights into the hyperparameters and experimental environment will be detailed in the following subsections.
5.3 Baselines
We compare our method with the following baselines, evaluated on all datasets and $F_{LLM}$ models:
SBERT-cosine: in this baseline, we employ a pre-trained SBERT model Reimers and Gurevych (2019) to embed both the query and the generated answer. We then calculate the cosine similarity between them and predict "hallucination" if the similarity falls below a threshold $SBERT_{\tau}$. The threshold was determined by using the same process described in Sec.5.2, this time with SBERT embeddings.
ADA-cosine: similar to SBERT-cosine but employs the recent openAI model ada002. The value of $\tau$ used here is consistent with the one in Sec.5.2.
5.4 The hallucination rates
We evaluate InterrogateLLM on answers generated by three recent LLMs for each of the datasets and tasks described above. The LLMs we evaluate are: GPT-3 Brown et al. (2020b) and Llama-2 Touvron et al. (2023b) (7b and 13b chat models). Interestingly, in Tab. 1 we report the hallucination rates in the generated answers of the three models across the different datasets. Notably, GPT-3 exhibits a lower hallucination rate across all datasets and tasks, compared to the Llama-2 models.
5.5 Hallucination detection results
Binary predictions (hallucinations or not) are compared to the ground truth test labels of each dataset. For each dataset and task, we employ InterrogateLLM with four different LLM choices for the backward step: GPT-3, Llama-2 (7B), and Llama-2 (13B), either individually or as ensembles of all three models. In Tab. 2, we report the area under the curve (AUC) of the receiver operating characteristic, and the balanced accuracy (B-ACC) metrics.
As can be seen in the table, the different variants of our method improve upon all the baselines by a sizeable margin. Within the context of the InterrogateLLM method’s different variants, our findings highlight GPT-3 as a superior $B_{LLM}$ compared to both Llama-2 variants across both the Movies and Books datasets. Notably, in cases where the $B_{LLM}$ performance with Llama-2 variants aligns closely with that of GPT-3, the ensemble approach proves highly advantageous. This implies that the combination of high-performing models further enhances the overall accuracy of the technique. Furthermore, when the ensemble attains performance levels similar to the GPT-3 backward model, we find that the Llama-2 variants exhibit weaker individual performance. In such instances, the ensemble effectively mitigates the weaknesses inherent in the Llama-2 models, resulting in an overall performance that closely aligns with that of the best single model. Additional insights into limitations are provided in Sec.6.
k=1 k=2 k=3 k=4 k=5
$F_{LLM}$ $B_{LLM}$   AUC    B-ACC     AUC    B-ACC     AUC    B-ACC     AUC    B-ACC     AUC    B-ACC
GPT3 GPT3 0.755 0.710 0.773 0.722 0.782 0.719 0.786 0.720 0.790 0.721
Llama-2 (7B) 0.701 0.633 0.721 0.641 0.727 0.635 0.732 0.638 0.734 0.631
Llama-2 (13B) 0.756 0.688 0.772 0.696 0.779 0.698 0.783 0.696 0.787 0.697
Ensemble 0.796 0.690 0.803 0.694 0.811 0.694 0.814 0.695 0.815 0.695
Llama-2 (7B) GPT3 0.775 0.774 0.786 0.778 0.788 0.776 0.794 0.782 0.798 0.780
Llama-2 (7B) 0.798 0.754 0.815 0.766 0.825 0.757 0.831 0.760 0.830 0.766
Llama-2 (13B) 0.810 0.782 0.824 0.778 0.828 0.780 0.836 0.781 0.838 0.783
Ensemble 0.840 0.786 0.850 0.787 0.852 0.790 0.853 0.792 0.853 0.795
Llama-2 (13B) GPT3 0.775 0.752 0.799 0.754 0.808 0.762 0.815 0.761 0.819 0.760
Llama-2 (7B) 0.757 0.704 0.763 0.710 0.764 0.701 0.767 0.702 0.769 0.699
Llama-2 (13B) 0.770 0.729 0.779 0.731 0.786 0.732 0.789 0.736 0.790 0.734
Ensemble 0.819 0.754 0.821 0.758 0.823 0.758 0.823 0.759 0.824 0.755
Table 3: Results for the Movies dataset. Results are reported for different k values, ranging from 1 to 5, average score. The highest AUC and B-ACC values for each row are presented in bold.
Figure 2: The average AUC and B-Acc scores across Movies, Books, and GCI datasets, per different K values (1-5).
Same temp Variable temp
$F_{LLM}$ $B_{LLM}$   AUC    B-ACC     AUC    B-ACC
GPT3 GPT3 0.790 0.721 0.817 0.739
Llama-2 (7B) 0.734 0.631 0.751 0.639
Llama-2 (13B) 0.787 0.697 0.789 0.695
Ensemble 0.815 0.695 0.818 0.699
Llama-2 (7B) GPT3 0.798 0.780 0.824 0.786
Llama-2 (7B) 0.830 0.766 0.823 0.750
Llama-2 (13B) 0.838 0.783 0.828 0.775
Ensemble 0.853 0.795 0.874 0.813
Llama-2 (13B) GPT3 0.819 0.760 0.806 0.753
Llama-2 (7B) 0.769 0.699 0.788 0.706
Llama-2 (13B) 0.790 0.734 0.801 0.746
Ensemble 0.824 0.755 0.842 0.773
Average 0.803 0.734 0.813 0.739
Table 4: Results for Movies dataset with same temperature and variable temperature.
5.6 Ablation and hyper-parameter analysis
We conduct an ablation study to examine the impact of the multiple $K$ backward processes performed in InterrogateLLM (Alg.1 line 4), the effectiveness of the variable temperature values (Eq.(8)), and the importance of the average function in Eq.5.
Variable $K$ values
The performance of InterrogateLLM with various values of $K$ is evaluated on the Movies, Books, and GCI datasets, and the results are reported in Tab. 3, and Tab. 9, 10 from the supplementary, respectively. Specifically, in this study, we evaluate InterrogateLLM with $K$ taking values in the range $[1,...,5]$ (higher $K$ values can be considered at the expense of more compute power). The tables reveal that utilizing $K>1$ in the backward step is crucial in all three experiments. Notably, the best results are consistently obtained with higher $K$ value. Therefore, we hypothesize that increasing the value of $K$ could potentially enhance the results, albeit at the expense of additional computational resources.
Fig. 2 depicts the enhancements araise from the different values of $K$, shown for each dataset seperately, and reported with both AUC and B-ACC metrics. Each data point represents the average result across all three forward LLMs along with all their corresponding backward LLMs (i.e. the average of each column in tables 3,9 and 10). As can be seen, the data reveals a consistent trend wherein the cumulative improvements exhibit a proportional relationship with the size of $K$.
Variable temperatures
We extend our investigation to varying temperatures for the backward process. For each index $i\in\text{range}(K)$, the InterrogateLLM method utilizes a variable temperature $T_{i}$ as defined in Eq.(8). This adjustment in temperature aimed to augment the creativity and stochastic aspects of the backward model $B_{LLM}$ throughout the query reconstruction process, fostering the generation of a more diverse set of reconstructed queries. Tab. 4, and Tab. 6, 7 from the supplementary Sec.A.3, present the results of InterrogateLLM with $K=5$, when using the same temperature through all the backward processes versus using variable temperatures, as proposed in InterrogateLLM . As can be seen, the variable temperature improves the results across most experiments in the Movies dataests, while yielding on par performance in the Books and GCI datasets (see Tab.6, 7). We hypothesize that the introduction of variable temperature, generating reconstructions with diverse levels of creativity, can be particularly helpful in mitigating instances of mode "collapse", in which certain backward models consistently generate identical reconstructions. In such cases, the incorporation of variable temperature becomes more important. The proposed method utilizes a diverse range of reconstructions. When the vast majority of these diverse reconstructions closely align with the original query, it signifies robust backward processes, better reflecting a non-hallucinated answer, and consequently leading to improved accuracy scores. More ablative experiments related to the choices made in Eq. 5 can be found in the sup. Sec.A.1.
6 Limitations
Throughout our study, we encountered several noteworthy limitations: (1) Source and Target Domain with Many-to-One Mapping: Generated answers associated with multiple different queries pose challenges in verification with InterrogateLLM . The backward process can reconstruct a diverse set of query candidates, deviating from the original query. (2) Hallucinating Back and Forth: Instances were observed where a single backward process by the same LLM model, which hallucinated an answer, could reconstruct the same query. This severe hallucination indicates a symmetric mapping between a query and a hallucinated answer, implying hallucinations in both directions. We observed a mitigation of this issue when employing an ensemble of models. (3) Detecting Hallucinations in Semi-Truth Answers: Identifying hallucinations in semi-truth answers proves more challenging. In some cases, the model only hallucinated a small portion of the answer (e.g., generating an entire cast of a movie with one additional actor not part of the movie). InterrogateLLM was able to recover the original movie, failing to detect the low-severity hallucination within the answer.
7 Conclusion
In this paper, we investigate the pressing issue of hallucinations in large language models.We introduced InterrogateLLM , a novel method designed for detecting hallucinations in few-shot settings. Our work contributes to the ongoing dialogue on the responsible use of AI-powered language models, offering a method that contributes to the reliability of LLM in diverse real-world applications.
As a future work, we would like to extend the method to Retrieval Augmented Generation settings, where a query is provided with a retrieved-context, and the task is to generate an answer based on the provided information in the context.
Supplementary Appendices
Appendix A More results
A.1 The Average-Max Analysis
For each query, we obtain a list of size K containing cosine similarity scores, representing the similarity between the generated query and the original query. To ascertain the closeness of the generated query to the original, we employ two strategies: the maximum (Max) and the average (Average) of the K cosine similarity scores. Notably, in the majority of cases, the average outperformed the maximum, indicating a more robust measure for query similarity. We attribute this observation to the sensitivity of the maximum strategy to noisy cosine similarity scores, which tend to be prevalent in certain instances. Our analysis suggests that the average is a more reliable indicator of query similarity, providing a stable and effective metric for the evaluation of our reconstruction approach. Full results with avg/max are presented in tables 8,9 and 10.
A.2 More results for variable $K$ values
Tab. 9 and 10 depicts the performance of InterrogateLLM on the books and GCI datasets, with variable $K$ values. As can be seen, higher values of $K$ result in enhanced accuracy.
A.3 More variable temperature results
Tab. 6 and 7 depicts the performance of the InterrogateLLM method with and without variable temperature values.
Appendix B More information about the datasets
B.1 Movies Dataset Query-Answer Examples
We provide query-answer examples from the Movies dataset, where each query seeks information about the cast of a specific movie:
  • Query-1: "What actors played in the 1995 movie Jumanji?"
  • Answer-1: "The main cast included Robin Williams, Jonathan Hyde, Kirsten Dunst, Bradley Pierce, Bonnie Hunt, Bebe Neuwirth, David Alan Grier, Patricia Clarkson, Adam Hann-Byrd, Laura Bell Bundy, James Handy, Gillian Barber, Brandon Obray, Cyrus Thiedeke, Gary Joseph Thorup, Leonard Zola, Lloyd Berry,Malcolm Stewart, Annabel Kershaw, Darryl Henriques, Robyn Driscoll, Peter Bryant, Sarah Gilson, Florica Vlad, June Lion, Brenda Lockmuller."
  • Query-2: "What actors played in the 2011 movie Kingdom Come?"
  • Answer-2: "Selma Blair, Edward Burns, Bruce Campbell, Lizzy Caplan, Seymour Cassel, Don Cheadle, Joanne Cook, Rachael Leigh Cook, Tom Cook, Alan Cumming,Tom DiCillo, Drake Doremus."
  • Query-3: "What actors played in the 2009 movie Inglourious Basterds?"
  • Answer-3: "The actors are Brad Pitt, Diane Kruger, Eli Roth, Mélanie Laurent, Christoph Waltz, Michael Fassbender, Daniel Brühl, Til Schweiger, Gedeon Burkhard, Jacky Ido, B.J. Novak, Omar Doom."
B.2 Books Dataset Query-Answer Examples
Query-Answer examples from the Books dataset, where queries inquire about the author and publication year of specific books:
  • Query-1: "Who is the author of the book Classical Mythology, and what year was it published?"
  • Answer-1: "The author is Mark P. O. Morford, and it was published in 2002."
  • Query-2: "Who is the author of the book Decision in Normandy, and what year was it published?"
  • Answer-2: "The author is Carlo D’Este, and it was published in 1991."
  • Query-3: "Who is the author of the book Clara Callan, what year was it published?"
  • Answer-3: "The author is Richard Bruce Wright, and it was published in 2001 by HarperFlamingo Canada."
B.3 GCI Dataset Query-Answer Examples
Query-Answer examples from the GCI dataset, where each query asks about the capital of a specific country:
  • Query-1: "What is the capital of France?"
  • Answer-1: "The capital is Paris."
  • Query-2: "What is the capital of Japan?"
  • Answer-2: "The capital is Tokyo."
  • Query-3: "What is the capital of Australia?"
  • Answer-3: "The capital is Canberra."
Same temp Variable temp
$F_{LLM}$ $B_{LLM}$   AUC    B-ACC     AUC    B-ACC
GPT3 GPT3 (avg) 0.790 0.721 0.817 0.739
GPT3 (max) 0.768 0.730 0.787 0.752
Llama-2 (7B) (avg) 0.734 0.631 0.751 0.639
Llama-2 (7B) (max) 0.721 0.669 0.726 0.690
Llama-2 (13B) (avg) 0.787 0.697 0.789 0.695
Llama-2 (13B) (max) 0.766 0.725 0.772 0.732
Ensemble (avg) 0.815 0.695 0.818 0.699
Ensemble (max) 0.786 0.741 0.798 0.756
Llama-2 (7B) GPT3 (avg) 0.798 0.780 0.824 0.786
GPT3 (max) 0.758 0.765 0.776 0.768
Llama-2 (7B) (avg) 0.830 0.766 0.823 0.750
Llama-2 (7B) (max) 0.814 0.781 0.808 0.773
Llama-2 (13B) (avg) 0.838 0.783 0.828 0.775
Llama-2 (13B) (max) 0.821 0.791 0.802 0.780
Ensemble (avg) 0.853 0.795 0.874 0.813
Ensemble (max) 0.802 0.765 0.810 0.772
Llama-2 (13B) GPT3 (avg) 0.819 0.760 0.806 0.753
GPT3 (max) 0.777 0.755 0.769 0.748
Llama-2 (7B) (avg) 0.769 0.699 0.788 0.706
Llama-2 (7B) (max) 0.757 0.728 0.772 0.738
Llama-2 (13B) (avg) 0.790 0.734 0.801 0.746
Llama-2 (13B) (max) 0.770 0.739 0.777 0.748
Ensemble (avg) 0.824 0.755 0.842 0.773
Ensemble (max) 0.763 0.733 0.792 0.756
Average (avg) 0.803 0.734 0.813 0.739
Average (max) 0.775 0.743 0.782 0.751
Table 5: Results for Movies dataset with same temperature and variable temperature.
Same temp Variable temp
$F_{LLM}$ $B_{LLM}$   AUC    B-ACC     AUC    B-ACC
GPT3 GPT3 (avg) 0.698 0.675 0.709 0.673
GPT3 (max) 0.685 0.670 0.694 0.667
Llama-2 (7B) (avg) 0.640 0.616 0.646 0.616
Llama-2 (7B) (max) 0.615 0.619 0.632 0.625
Llama-2 (13B) (avg) 0.675 0.642 0.684 0.640
Llama-2 (13B) (max) 0.656 0.643 0.669 0.648
Ensemble (avg) 0.707 0.656 0.710 0.656
Ensemble (max) 0.707 0.669 0.719 0.681
Llama-2 (7B) GPT3 (avg) 0.821 0.777 0.828 0.787
GPT3 (max) 0.811 0.780 0.815 0.784
Llama-2 (7B) (avg) 0.761 0.707 0.761 0.707
Llama-2 (7B) (max) 0.744 0.718 0.752 0.725
Llama-2 (13B) (avg) 0.794 0.730 0.795 0.734
Llama-2 (13B) (max) 0.783 0.745 0.785 0.752
Ensemble (avg) 0.824 0.769 0.822 0.761
Ensemble (max) 0.831 0.793 0.827 0.783
Llama-2 (13B) GPT3 (avg) 0.799 0.757 0.804 0.754
GPT3 (max) 0.792 0.758 0.797 0.763
Llama-2 (7B) (avg) 0.743 0.686 0.742 0.679
Llama-2 (7B) (max) 0.722 0.696 0.731 0.707
Llama-2 (13B) (avg) 0.771 0.707 0.771 0.709
Llama-2 (13B) (max) 0.754 0.714 0.759 0.724
Ensemble (avg) 0.802 0.739 0.807 0.733
Ensemble (max) 0.808 0.765 0.817 0.774
Average (avg) 0.752 0.705 0.756 0.704
Average (max) 0.742 0.714 0.747 0.719
Table 6: Results for Books dataset with same temperature and variable temperature.
Same temp Variable temp
$F_{LLM}$ $B_{LLM}$   AUC    B-ACC     AUC    B-ACC
GPT3 GPT3 (avg) - 0.994 - 0.994
GPT3 (max) - 0.994 - 0.994
Llama-2 (7B) (avg) - 0.983 - 0.983
Llama-2 (7B) (max) - 0.983 - 0.983
Llama-2 (13B) (avg) - 0.983 - 0.983
Llama-2 (13B) (max) - 0.983 - 0.983
Ensemble (avg) - 0.983 - 0.983
Ensemble (max) - 0.983 - 0.983
Llama-2 (7B) GPT3 (avg) 0.969 0.972 0.965 0.952
GPT3 (max) 0.968 0.972 0.964 0.952
Llama-2 (7B) (avg) 0.974 0.957 0.959 0.958
Llama-2 (7B) (max) 0.976 0.961 0.960 0.962
Llama-2 (13B) (avg) 0.977 0.959 0.969 0.960
Llama-2 (13B) (max) 0.977 0.959 0.971 0.967
Ensemble (avg) 0.963 0.951 0.951 0.948
Ensemble (max) 0.944 0.944 0.949 0.941
Llama-2 (13B) GPT3 (avg) 0.986 0.982 0.989 0.982
GPT3 (max) 0.971 0.978 0.983 0.977
Llama-2 (7B) (avg) 1.000 1.000 1.000 1.000
Llama-2 (7B) (max) 1.000 1.000 1.000 1.000
Llama-2 (13B) (avg) 1.000 0.991 0.995 0.991
Llama-2 (13B) (max) 0.998 0.978 0.983 0.986
Ensemble (avg) 1.000 0.991 0.992 0.964
Ensemble (max) 0.987 0.986 0.983 0.967
Average (avg) 0.983 0.974 0.977 0.974
Average (max) 0.977 0.976 0.974 0.974
Table 7: Results for GCI dataset with same temperature and variable temperature.
k=1 k=2 k=3 k=4 k=5
$F_{LLM}$ $B_{LLM}$   AUC    B-ACC     AUC    B-ACC     AUC    B-ACC     AUC    B-ACC     AUC    B-ACC
GPT3 GPT3 (max) 0.755 0.710 0.765 0.724 0.768 0.730 0.767 0.729 0.768 0.730
GPT3 (avg) 0.755 0.710 0.773 0.722 0.782 0.719 0.786 0.720 0.790 0.721
Llama-2 (7B) (max) 0.701 0.633 0.714 0.650 0.714 0.659 0.718 0.664 0.721 0.669
Llama-2 (7B) (avg) 0.701 0.633 0.721 0.641 0.727 0.635 0.732 0.638 0.734 0.631
Llama-2 (13B) (max) 0.756 0.688 0.761 0.707 0.764 0.715 0.765 0.721 0.766 0.725
Llama-2 (13B) (avg) 0.756 0.688 0.772 0.696 0.779 0.698 0.783 0.696 0.787 0.697
Ensemble (max) 0.782 0.736 0.778 0.742 0.785 0.744 0.787 0.745 0.786 0.741
Ensemble (avg) 0.796 0.690 0.803 0.694 0.811 0.694 0.814 0.695 0.815 0.695
Llama-2 (7B) GPT3 (max) 0.775 0.774 0.767 0.775 0.761 0.769 0.758 0.766 0.758 0.765
GPT3 (avg) 0.775 0.774 0.786 0.778 0.788 0.776 0.794 0.782 0.798 0.780
Llama-2 (7B) (max) 0.798 0.754 0.808 0.775 0.812 0.778 0.818 0.782 0.814 0.781
Llama-2 (7B) (avg) 0.798 0.754 0.815 0.766 0.825 0.757 0.831 0.760 0.830 0.766
Llama-2 (13B) (max) 0.810 0.782 0.812 0.781 0.814 0.785 0.822 0.791 0.821 0.791
Llama-2 (13B) (avg) 0.810 0.782 0.824 0.778 0.828 0.780 0.836 0.781 0.838 0.783
Ensemble (max) 0.810 0.782 0.810 0.776 0.810 0.773 0.809 0.770 0.802 0.765
Ensemble (avg) 0.840 0.786 0.850 0.787 0.852 0.790 0.853 0.792 0.853 0.795
Llama-2 (13B) GPT3 (max) 0.775 0.752 0.779 0.755 0.775 0.753 0.777 0.752 0.777 0.755
GPT3 (avg) 0.775 0.752 0.799 0.754 0.808 0.762 0.815 0.716 0.819 0.760
Llama-2 (7B) (max) 0.757 0.704 0.758 0.717 0.754 0.721 0.753 0.727 0.757 0.728
Llama-2 (7B) (avg) 0.757 0.704 0.763 0.710 0.764 0.701 0.767 0.702 0.769 0.699
Llama-2 (13B) (max) 0.770 0.729 0.770 0.732 0.770 0.735 0.770 0.736 0.770 0.739
Llama-2 (13B) (avg) 0.770 0.729 0.779 0.731 0.786 0.732 0.789 0.736 0.790 0.734
Ensemble (max) 0.793 0.765 0.777 0.751 0.774 0.743 0.766 0.739 0.763 0.733
Ensemble (avg) 0.819 0.754 0.821 0.758 0.823 0.758 0.823 0.759 0.824 0.755
Table 8: Results for the Movies dataset. Results are reported for different k values, ranging from 1 to 5, average and max scores.
k=1 k=2 k=3 k=4 k=5
$F_{LLM}$ $B_{LLM}$   AUC    B-ACC     AUC    B-ACC     AUC    B-ACC     AUC    B-ACC     AUC    B-ACC
GPT3 GPT3 (max) 0.680 0.657 0.683 0.671 0.681 0.669 0.682 0.669 0.685 0.670
GPT3 (avg) 0.680 0.657 0.691 0.673 0.692 0.676 0.695 0.681 0.698 0.675
Llama-2 (7B) (max) 0.626 0.606 0.621 0.615 0.613 0.614 0.610 0.616 0.609 0.617
Llama-2 (7B) (avg) 0.626 0.606 0.635 0.615 0.633 0.614 0.634 0.614 0.634 0.615
Llama-2 (13B) (max) 0.654 0.623 0.654 0.634 0.659 0.640 0.660 0.643 0.656 0.643
Llama-2 (13B) (avg) 0.654 0.623 0.665 0.633 0.670 0.637 0.673 0.638 0.675 0.642
Ensemble (max) 0.693 0.668 0.696 0.670 0.698 0.668 0.703 0.668 0.707 0.669
Ensemble (avg) 0.696 0.658 0.703 0.663 0.704 0.657 0.706 0.655 0.707 0.656
Llama-2 (7B) GPT3 (max) 0.795 0.757 0.804 0.771 0.804 0.773 0.809 0.777 0.811 0.780
GPT3 (avg) 0.795 0.757 0.811 0.772 0.815 0.773 0.820 0.774 0.821 0.777
Llama-2 (7B) (max) 0.737 0.686 0.744 0.704 0.746 0.712 0.743 0.714 0.744 0.718
Llama-2 (7B) (avg) 0.737 0.686 0.754 0.703 0.760 0.708 0.760 0.709 0.761 0.707
Llama-2 (13B) (max) 0.773 0.720 0.778 0.734 0.779 0.738 0.781 0.741 0.783 0.745
Llama-2 (13B) (avg) 0.773 0.720 0.785 0.729 0.791 0.732 0.793 0.731 0.794 0.730
Ensemble (max) 0.806 0.777 0.818 0.787 0.822 0.789 0.827 0.793 0.831 0.793
Ensemble (avg) 0.811 0.766 0.817 0.768 0.819 0.764 0.822 0.764 0.824 0.769
Llama-2 (13B) GPT3 (max) 0.776 0.733 0.782 0.745 0.783 0.748 0.788 0.755 0.792 0.758
GPT3 (avg) 0.776 0.733 0.789 0.750 0.794 0.755 0.797 0.754 0.799 0.757
Llama-2 (7B) (max) 0.716 0.674 0.721 0.688 0.724 0.695 0.722 0.695 0.722 0.696
Llama-2 (7B) (avg) 0.716 0.674 0.732 0.689 0.740 0.690 0.743 0.690 0.743 0.686
Llama-2 (13B) (max) 0.748 0.694 0.749 0.706 0.749 0.709 0.751 0.714 0.754 0.714
Llama-2 (13B) (avg) 0.748 0.694 0.761 0.707 0.764 0.703 0.769 0.707 0.771 0.707
Ensemble (max) 0.788 0.752 0.800 0.765 0.803 0.765 0.807 0.766 0.808 0.765
Ensemble (avg) 0.793 0.736 0.800 0.741 0.801 0.739 0.801 0.737 0.802 0.739
Table 9: Results for the Books dataset. Results are reported for different k values, ranging from 1 to 5, with average and max scoring.
k=1 k=2 k=3 k=4 k=5
$F_{LLM}$ $B_{LLM}$   AUC    B-ACC     AUC    B-ACC     AUC    B-ACC     AUC    B-ACC     AUC    B-ACC
GPT3 GPT3 (max) - 0.994 - 0.994 - 0.994 - 0.994 - 0.994
GPT3 (avg) - 0.994 - 0.994 - 0.994 - 0.994 - 0.994
Llama-2 (7B) (max) - 0.983 - 0.983 - 0.983 - 0.983 - 0.983
Llama-2 (7B) (avg) - 0.983 - 0.983 - 0.983 - 0.983 - 0.983
Llama-2 (13B) (max) - 0.983 - 0.983 - 0.983 - 0.983 - 0.983
Llama-2 (13B) (avg) - 0.983 - 0.983 - 0.983 - 0.983 - 0.983
Ensemble (max) - 0.983 - 0.983 - 0.983 - 0.983 - 0.983
Ensemble (avg) - 0.983 - 0.983 - 0.983 - 0.983 - 0.983
Llama-2 (7B) GPT3 (max) 0.969 0.968 0.969 0.968 0.969 0.972 0.968 0.972 0.968 0.972
GPT3 (avg) 0.969 0.968 0.969 0.968 0.969 0.972 0.970 0.972 0.969 0.972
Llama-2 (7B) (max) 0.975 0.957 0.976 0.961 0.976 0.961 0.976 0.961 0.976 0.961
Llama-2 (7B) (avg) 0.975 0.957 0.975 0.961 0.974 0.961 0.974 0.957 0.974 0.957
Llama-2 (13B) (max) 0.977 0.959 0.977 0.959 0.977 0.959 0.977 0.959 0.977 0.959
Llama-2 (13B) (avg) 0.977 0.959 0.977 0.959 0.977 0.959 0.977 0.959 0.977 0.959
Ensemble (max) 0.945 0.944 0.944 0.944 0.944 0.944 0.944 0.944 0.944 0.944
Ensemble (avg) 0.964 0.951 0.963 0.951 0.964 0.951 0.963 0.951 0.963 0.951
Llama-2 (13B) GPT3 (max) 0.980 0.982 0.980 0.982 0.980 0.982 0.980 0.982 0.971 0.978
GPT3 (avg) 0.980 0.982 0.980 0.982 0.980 0.982 0.986 0.982 0.986 0.982
Llama-2 (7B) (max) 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Llama-2 (7B) (avg) 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000 1.000
Llama-2 (13B) (max) 1.000 0.991 1.000 0.991 0.998 0.982 0.998 0.982 0.998 0.978
Llama-2 (13B) (avg) 1.000 0.991 1.000 0.991 1.000 0.991 1.000 0.991 1.000 0.991
Ensemble (max) 0.987 0.991 0.987 0.991 0.987 0.991 0.987 0.991 0.987 0.986
Ensemble (avg) 1.000 0.991 1.000 0.995 1.000 0.991 1.000 0.991 1.000 0.991
Table 10: Results for the GCI dataset. Results are reported for different k values, ranging between 1 to 5, average scoring.